OpenStudy (ashy98):

can some one help me with 4 questions??(:

OpenStudy (ashy98):

1.) If you were solving a system of equations and you came to a statement like 4 = 4, what do you know about the solution to the system? A.)The solution is (4, 4) B.) The solution is x = 4 and y = 4 C.) There is no solution D.) There are infinitely many solutions 2.) Choose the equivalent system of linear equations that will produce the same solution as the one given below. 4x - 2y = 6 2x + y = 5 A.) -4x - 2y = 10 -4y = 4 B.) -4x - 5y = -1 -7y = 5 C.) 3x + 2y = 6 7x = 12 D.) 4x + 2y = 10 8x = 16 3.) Determine which system below will produce infinitely many solutions. A.) -6x + 3y = 18 4x - 3y = 6 B.) 2x + 4y = 24 6x + 12y = 36 C.) 3x - y = 14 -9x + 3y = -42 D.) 5x + 2y = 13 -x + 4y = -6 4.) Solve the system of equations. x + 3y = -1 2x + 2y = 6 A.) (-4, 1) B.) (2, -1) C.) (4, -1) D.) (5, -2)

OpenStudy (ashy98):

@rishavraj

OpenStudy (ashy98):

@SithsAndGiggles @TheDream @whpalmer4

OpenStudy (triciaal):

many solutions

OpenStudy (ganpat):

For 1st question.. 4 = 4.. take everything on one side.. 4-4 = 0 0 = 0 So, No Solution

OpenStudy (ganpat):

orr.. infinite solution.. m not sure ( let me try )

OpenStudy (triciaal):

2) divide eqn 1 by 2 2x -y = 3

OpenStudy (ashy98):

for #4 i got (5,-2)

OpenStudy (ganpat):

4th one.. x + 3y = -1 2x + 2y = 6 Substitute these values in both the equations separately.. A.) (-4, 1) = ( -1 = -1) (-6 = 6) ---- No B.) (2, -1) = ( -1 = -1) (2 = 6) ----- No C.) (4, -1) = ( 1 = -1) ---- No D.) (5, -2) = ( -1 = -1) ( 6 = 6 ) ---- is the answer..

OpenStudy (triciaal):

add the 2 eqns 4x = 8 so x = 2 substitute to find y 2x + y = 5 y = 1 now find which choice is true when x = 2 and y = 1

OpenStudy (ashy98):

thats how i got my answer (: @Ganpat

OpenStudy (ganpat):

3.) Determine which system below will produce infinitely many solutions. A.) -6x + 3y = 18 4x - 3y = 6 B.) 2x + 4y = 24 6x + 12y = 36 C.) 3x - y = 14 -9x + 3y = -42 D.) 5x + 2y = 13 -x + 4y = -6 For infinitely many solution.. it has to be one and the same equation.. for example.. x + y = 1 and 2x + 2y = 2... this has infinitely many solutions. U can divide equation 1 & 2 for each of the option.. C) 1st equation is multiplied by -3... So C is the answer..

OpenStudy (ganpat):

2.) Choose the equivalent system of linear equations that will produce the same solution as the one given below. 4x - 2y = 6 2x + y = 5.. multiply by 2 4x + 2y = 10... eq 2 4x - 2y = 6...... eq 1 8x = 16.... addition x = 2 4(2) -2(y) = 6 y = 1 Substitute value of x & y in given options... and you will find your answer... D.) 4x + 2y = 10 8x = 16

OpenStudy (ganpat):

Of course, I answer is Infinite solutions... It represents a line.. ha haa